Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, aaron.jones, gherardini - 21:36 Tuesday 13 May 2025 (66765) Print this report
ITF recovery

After the recovery of the West Arm and the end of Ncal operations, we started the recovery of the ITF without WE inertial damping.

The recovery proceeded smoothly until the lock of the DRMI, but we couldn't continue because the galvo loops would not stay closed. The DRMI lock itself is unstable, and does not last more than a few minutes, because the alignment loops are not working properly.

The beam on B1p camera looks reasonably well centered, so it might not be related to a bench misalignment.

It is not clear whether this is related to the absence of inertial damping on the WE, we contacted Diego and, as instructed, we leave the ITF as is.

The recovery will be continued tomorrow.

Comments to this report:
bersanetti - 21:55 Tuesday 13 May 2025 (66766) Print this report

What we see looks similar to something that happened recently (#66258), and some interplay between SR and SDB1 alignment was the cause, I don't know if that is the case also now.

mwas - 8:34 Wednesday 14 May 2025 (66768) Print this report

This is a good reminder.

 Figure 1. shows that SDB1 has changed orientation compared to where it was a few weeks ago, and during the recovery it has moved further away.

This morning I have disabled the floating set points (the bench learning and remembering on its own what are the set points), and put back the set points to the average values of the bench orientation in TX and TY from beginning of May.

We will see if helps, the interferometer locked on DRMI for 20 seconds, and the B1p quadrants galvo close during that time.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×